OpenStudy (anonymous):

Please please help! I keep getting horribly wrong asnwers! x - 3y + 4z = -9 2x + y + 3z = 6 x - 4y + 3z = -9

OpenStudy (anonymous):

A) x - 3y + 4z = -9 B) 2x + y + 3z = 6 C) x - 4y + 3z = -9 A+B+C = ? say...

OpenStudy (anonymous):

\[\begin{cases}x-3y+4z=-9\\2x+y+3z=6\\x-4y+3z=-9\end{cases}\] Solving for \(x\) in the third equation: \[x=4y-3z-9\] Substituting this into both the first and second equation (thus eliminating \(x\): \[\begin{cases}(4y-3z-9)-3y+4z=-9\\2(4y-3z-9)+y+3z=6\end{cases}~~\Rightarrow~~\begin{cases}y+z=0\\9y-3z=24\end{cases}\] Solving for \(y\) in the first equation yields \(y=-z\). Plug this into the second equation: \[9(-z)-3z=24~~\Rightarrow~~-12z=24~~\Rightarrow~~z=-2\] \[z=-2~\Rightarrow~y=2~\Rightarrow~x=5\]